# On an entropy inequality for quadratic forms and applications

### v2

@@ 2.3 Discretized entropy and properties
$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots,Y_{l})$
$\quad=H(D_{n}(X_{l})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
@@
In [16], Ruzsa studied some entropy inequalities involving sums of random variables. In the following, more Plünnecke-Ruzsa type inequalities for discretized entropy are presented.
@@
Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist a non-negative integers $d=d(k,l)$ such that
$H_{n}(kX+lY)\leq dH_{n}(X+Y)-(d-1)(H_{n}(X)-H_{n}(Y))+O_{k,l}(1)$
@@
We first prove that
$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)$
$\qquad-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+(d_{1}+d_{2}+1)I_{n}(X,Y)+O_{k,l}(1)$ (14)
for some non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$.
The proof is trivial when $k=l=0$. When $l=0$ and $k\neq 0$, we have
$H_{n}(kX)=H_{n}(X)+O_{k}(1)\leq H_{n}(X+Y)+O_{k}(1)$
by (13). Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
First, let us prove the case that $k,l\in\mathbb{Z}^{\ast}$. We apply double induction on $|k|$ and $|l|$. When $|k|=|l|=1$, we claim that
$H_{n}(kX+lY)\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (15)
Indeed, when $k=l=1$, one can deduce (15) from (13). When $k=1$ and l=1l=-1, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed. Then Corollary 2.20 gives
$H_{n}(X-Y)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
By Corollary 2.20 again, we have
$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1).$ (16)
Then (15) follows. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
Now let us assume that the conclusion has been proved for some $k$ and $l$. Without loss of generality, we assume that $k>0$. We aim to prove the conclusion for $k+1$ and $l$. Lemma 2.17 gives us that
$H_{n}((k+1)X+lY+Z)+H_{n}(X,Y,Z)\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)+O_{k,l}(1),$
where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining (13), we obtain that
$H_{n}((k+1)X+lY)\leq H_{n}((k+1)X+lY+Z)+O_{k,l}(1)$
$\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)-H_{n}(X,Y,Z)+O_{k,l}(1)$
$\leq H_{n}(Z+lY)+H_{n}(X+Z)+H_{n}(kX+lY)-H_{n}(Y)-H_{n}(Z)+O_{k,l}(1).$
Combining the fact
$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1)$
and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
$d_{1}(k+1,l)=d_{1}(1,l)+d_{1}(k,l)+1,\quad d_{2}(k,l)=d_{2}(1,l)+d_{2}(k,l)+2.$
By symmetry, we also have the induction on $l$, and the proof of (14) is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
@@
Second, when $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It follows that
$H_{n}(kX+lY)$ $=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$
$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k^{\prime},l^{\prime},t}(1)$
for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. Then (14) follows.
Next, in view of (15), we can add multiple scalars of $H_{n}(X+Y)-H_{n}(X)$ and $H_{n}(X+Y)-H_{n}(Y)$ on the right-hand side of (14) to get
$H_{n}(kX+lY)\leq(2d_{1}+2d_{2}+1)H_{n}(X+Y)-(2d_{1}+2d_{2})\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1).$
The lemma then follows by taking $d=2d_{1}+2d_{2}+1$. ∎
@@
Lemma 2.21 implies the following statement: For given $k,l\in\mathbb{Q}$, one has
$H_{n}(kX+lY)-H_{n}(X)\ll_{k,l}(H_{n}(X+Y)-H_{n}(X))+O_{k,l}(1)$
@@
for any $n\in\mathbb{N}$ and any i.i.d. random variables $X$, $Y$ with bounded range. However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail. In Example 7.8, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.18,
@@
Using Lemmas 2.15 and 2.17, we have that
@@
4 Proof of Theorems 1.2-1.4 with $s>1/2$
@@ 4.1 Diagonal forms with independent random variables
$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (35)
@@
For the third term, we use Lemma 2.20 repeatedly, and combine Lemma 2.15 and formula (36), to get
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
$\leq H_{n}(p(X_{1})+p(X_{2})+q(Y_{1}))+H_{n}(p(X_{1})+q(Y_{1})+q(Y_{2}))-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
$\leq H_{n}(p(X_{1})+q(Y_{1}))+H_{n}(p(X_{2})+q(Y_{1}))-H_{n}(q(Y_{1}))$
$\qquad+H_{n}(p(X_{1})+q(Y_{2}))+H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))$
$\qquad-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (39)
@@
Compared with (39), the second entropy can be bounded by
$H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))\leq 3H_{n}(p(X_{1})-q(Y_{1}))-H_{n}(X_{1})-H_{n}(Y_{1})+O(1).$
The term $H_{n}(p(X_{1})-q(Y_{1}))$ can be bounded by using (15), which gives
$H_{n}(p(X_{1})-q(Y_{1}))\leq 3H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))-H_{n}(q(Y_{1}))+O(1)$
$\quad=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
@@ 4.2 Diagonal forms with dependent random variables
Theorem 1.4’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
@@
Next, we prove Theorem 1.3 in the following form.
Theorem 1.3’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
(i) For all $n\in\mathbb{N}$,

### v3

@@ 2.3 Discretized entropy and properties
Example 2.17.
Let $m\geq 1$. Suppose that $A$ is an $m\times m$ real matrix and $X,X_{1},X_{2}$ are multivariate real random variables of dimension $m$. Note that $|A(X_{1}-X_{2})|\leq\|A\|\cdot|X_{1}-X_{2}|$, where $\|A\|$ is the operator norm of $A$. Lemma 2.15 shows that $H_{n}(AX)\leq H_{n}(X)+O_{A}(1)$. In particular, if $A$ is invertible, then $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
@@
$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots+Y_{l})$
$\quad=H(D_{n}(X_{1})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
@@
In [16], Ruzsa studied some entropy inequalities that involving sums of random variables. In the following, more inequalities for discretized entropy are presented.
@@
Without independence, the follow lemma can be used as an alternative.
@@
Let $X,Y$ and $Z$ be real random variables that are not necessarily independent. Then, for all $n\in\mathbb{N}$, we have
$H_{n}(X+Y)\leq H_{n}(X+Y+Z)+H_{n}(X,Y)+H_{n}(Z)-H_{n}(X,Y,Z)+O(1),$
and
$H_{n}(X+Y)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-2H_{n}(X,Y,Z)$
$\quad+H_{n}(X,Y)+H_{n}(X)+H_{n}(Y)+H_{n}(Z)+O(1).$
Proof.
Lemma 2.18 shows that
$H_{n}(X+Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Y+Z,Z)+H_{n}(X,Y)+O(1).$
Combining (14), The lemma then follows. ∎
An important task of this subsection is to prove the following theorem on the relationship between $H_{n}(kX+lY)$ and $H_{n}(X+Y)$.
Theorem 2.23.
Let $k,l\in\mathbb{Q}^{\ast}$, and $X$ and $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)$
$\qquad\qquad\qquad+d_{2}\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+O_{k,l}(1)$
@@
The proof is divided into several lemmas as below.
Lemma 2.24.
Let $X$ and $Y$ be real random variables that are not necessarily independent. Then, for all $n\in\mathbb{N}$,
$H_{n}(X-Y)$ $\leq 3H_{n}(X+Y)+4H_{n}(X)+4H_{n}(Y)-5H_{n}(X,Y)+O(1).$
@@
For a given $n\in\mathbb{N}$, let $X^{\prime}$ be a random variable that is conditionally i.i.d. with $X$ given $D_{n}(Y)$. Then $D_{n}(X)$ and $D_{n}(X^{\prime})$ are also conditionally i.i.d. given $D_{n}(Y)$, and $(D_{n}(X),D_{n}(Y))\equiv(D_{n}(X^{\prime}),D_{n}(Y))$. By Lemma 2.8,
$H_{n}(X,X^{\prime},Y)=2H_{n}(X,Y)-H_{n}(Y).$ (16)
Recall Lemma 2.20. By (15), one has
$H_{n}(X-Y)+H_{n}(X,X^{\prime},Y)\leq H_{n}(X+X^{\prime})+H_{n}(Y+X^{\prime})+H_{n}(X,Y)+O(1).$ (17)
In view of the second inequality in Lemma 2.22, we get
$\begin{split}H_{n}(X+X^{\prime})&\leq H_{n}(X+Y)+H_{n}(X^{\prime}+Y)-2H_{n}(X,X^{\prime},Y)\\ &\quad+H_{n}(X,X^{\prime})+H_{n}(X)+H_{n}(X^{\prime})+H_{n}(Y)+O(1).\end{split}$ (18)
Combining (16)-(18), the lemma then follows. ∎
Remark 2.25.
When $X$ and $Y$ are independent, one has
$H_{n}(X-Y)$ $\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$
for all $n\in\mathbb{N}$.
For simplicity, we will denote the discretized mutual information by
$\mathcal{I}_{n}(X,Y)=H_{n}(X)+H_{n}(Y)-H_{n}(X,Y).$
Lemma 2.26.
Let $k\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
$H_{n}(kX+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O_{k}(1).$
for all $n\in\mathbb{N}$.
Proof.
For $k=1$, the conclusion holds for $d_{1}=1$ and $d_{2}=0$. For k=1k=-1, Lemma 2.24 shows that we can take $d_{1}=3$ and $d_{2}=5$. Now we apply induction on $|k|$. Assume that the conclusion has been proved for $k$ with $|k|=m$, where $m\geq 1$. We consider the situation $k=\pm(m+1)$.
Let $X^{\prime}$ be as in the proof of Lemma 2.24. By the first inequality in Lemma 2.22,
$\begin{split}&H_{n}\big{(}\pm(m+1)X+Y\big{)}\leq H_{n}\big{(}\pm(m+1)X+Y\pm mX^{\prime}\big{)}\\ &\qquad\qquad\qquad+H_{n}(X,Y)+H_{n}(X^{\prime})-H_{n}(X,X^{\prime},Y)+O_{m}(1).\end{split}$ (19)
Note that
$\pm(m+1)X+Y\pm mX^{\prime}=\pm(m+1)(X\pm Y)\pm m(X^{\prime}\mp Y).$
Applying Lemma 2.18 yields
$H_{n}\big{(}\pm(m+1)X+Y\pm mX^{\prime}\big{)}+H_{n}(X,X^{\prime},Y)$
$\qquad\leq H_{n}(\pm mX^{\prime}+Y,X)+H_{n}(X\pm Y,X^{\prime}\mp Y)+O(1)$
$\qquad\leq H_{n}(\pm mX+Y)+H_{n}(X+Y)+H_{n}(X-Y)+H_{n}(X)+O(1).$ (20)
Combining Lemma 2.24, (16), (19) and (20), it follows that
$H_{n}\big{(}\pm(m+1)X+Y\big{)}\leq H_{n}(\pm mX+Y)+4H_{n}(X+Y)+6H_{n}(X)+6H_{n}(Y)-8H_{n}(X,Y).$
Therefore, the conclusion follows with
$d_{1}(\pm(m+1))=d_{1}(\pm m)+4,\qquad d_{1}(\pm(m+1))=d_{1}(\pm m)+8.$
Lemma 2.27.
Let $k,l\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O(1).$
for all $n\in\mathbb{N}$.
Proof.
Applying Lemma 2.26 with the random variable $(X,lY)$, we obtain that
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
$H_{n}(lY+X)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
$\qquad\leq d_{1}(l,1)\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(l,1)\mathcal{I}_{n}(X,Y)+O_{l}(1).$
Hence, the conclusion follows by taking
$d_{1}(k,l)=d_{1}(k,1)d_{1}(l,1),\qquad d_{2}(k,l)=d_{1}(k,1)d_{2}(l,1)+d_{2}(k,1).$
Proof of Theorem 2.23.
@@
Since $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It satisfies that $H_{n}(kX+lY)=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$. The conclusion follows from Lemma 2.27. ∎
Remark 2.28.
@@
Theorem 2.23 implies the following statement: For all given $k,l\in\mathbb{Q}$, if $X$ and $Y$ are independent random variables, then
$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\ll_{k,l}H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1)$
for any $n\in\mathbb{N}$. Here, the case that $k=0$ or $l=0$ can be proved by using (13).
@@
However, when $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail. In Example 7.7, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@
Lemma 2.29.
Let $X,Y$ be independent random variables, and $f,g$ be weakly $\kappa$-bi-Lipschitz real-valued functions. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. Then
$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa),$
$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa),$
$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa).$
Proof.
Note that $f(X_{1}),f(X_{2}),g(Y_{1}),g(Y_{2})$ are independent, and $(X_{1},Y_{1})\equiv(X_{2},Y_{2})\equiv(X,Y)$. By using Lemma 2.21 repeatedly, combining Lemma 2.15, we get
$H_{n}(f(X_{1})\pm f(X_{2})+g(Y_{1})\pm g(Y_{2}))$
$\leq H_{n}(f(X_{1})\pm f(X_{2})+g(Y_{1}))+H_{n}(f(X_{1})+g(Y_{1})\pm g(Y_{2}))-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
$\leq H_{n}(f(X_{1})+g(Y_{1}))+H_{n}(\pm f(X_{2})+g(Y_{1}))-H_{n}(g(Y_{1}))$
$\qquad+H_{n}(f(X_{1})+g(Y_{1}))+H_{n}(f(X_{1})\pm g(Y_{2}))-H_{n}(f(X_{1}))$
$\qquad-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
By Remark 2.25, it satisfies that
$H_{n}(f(X)-g(Y))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(f(X))-H_{n}(q(Y))+O(1)$
$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
The above formulae still hold if $g$ is replaced by g-g. The lemma then follows from calculation. ∎
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.19,
@@
Using Lemmas 2.15 and 2.18, we have that
@@
4 Proof of Theorems 1.21.4 with $s>1/2$
@@ 4.1 Diagonal forms with independent random variables
$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (39)
@@
For the third term, Lemma 2.29 shows that
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 3H_{n}(p(X)+q(Y))$
$\qquad\qquad\qquad-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (43)
@@
Compared with (43), Lemma 2.29 shows that
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 9H_{n}(p(X)+q(Y))$
$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
@@ 4.2 Diagonal forms with dependent random variables
Theorem 1.2’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C,c>0$. There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
@@
4.3 The general forms
Any non-degnerate rational quadratic from $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ can be diagonalized by rational linear change of variables, i.e.,
for some $\alpha,\beta\in\mathbb{Q}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
One can verify that $\phi(x,y)$ is diagonal if and only if exactly two of the coefficients $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ are zero. For non-diagonal case, at most one of $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ may be zero. In the following, we will show that, if exactly one of these numbers is zero, then we can transform $\phi(x,y)$ into another form like (56) with all corresponding coefficients non-zero. Now we assume without loss of generality that $\xi_{11},\xi_{21},\xi_{22}\neq 0$ and $\xi_{21}=0$. By dividing $\xi_{11}$ and $\xi_{12}$ in the two brackets, respectively, we may rewrite $\phi(x,y)=\alpha^{\prime}x^{2}+\beta^{\prime}(x+d^{\prime}y)^{2}$, with $\alpha^{\prime},\beta^{\prime},d^{\prime}\in\mathbb{Q}^{\ast}$. Now, let us take some $a,c\in\mathbb{Q}^{\ast}$ satisfying $a\neq c$, $a\neq 1/d^{\prime}$, $c\neq 1/d^{\prime}$ and
$(d^{\prime}a-1)(d^{\prime}c-1)=-\alpha^{\prime}/\beta^{\prime}.$
One can check that
$\phi(x,y)=\frac{\beta^{\prime}d^{\prime}(1-cd^{\prime})}{a-c}(ax+y)^{2}+\frac{\beta^{\prime}d^{\prime}(ad^{\prime}-1)}{a-c}(cx+y)^{2}.$
All coefficient in above formula are non-zero rational numbers. Furthermore, we can assume in (56) without loss of generality that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$, because non-zero constants can be multiplied inside both brackets. In all situations, all the coefficients depend on $a_{1},a_{2},a_{3}$.
Now we prove Theorem 1.3 in the form of the following two theorems, which hold for more general forms.
Theorem 1.3’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then